3.10 A rotating thin steel cylinder is replaced by a rotating thin aluminium cylinder

of the same mass, length and wall thickness. The stress due to rotation in

the aluminium cylinder is to be 25 % of that in the steel cylinder. Calculate the

allowable speed of the aluminium cylinder if the steel cylinder rotates at 600

r/min (relative density of steel = 7,8; relative density of aluminium = 2,9)


1
Expert's answer
2021-03-23T08:04:40-0400

The force that provides the stress depends on the tangential speed of rotation at the walls of the cylinder:


"f=k\rho\frac{v^2}{R}=k\rho\omega^2R,\\\space\\
f_a=0.25f_s,\\\space\\
\rho_a\omega_a^2R=0.25\rho_s\omega_s^2R,\\\space\\
\omega_a=0.25\omega_s\sqrt{\frac{\rho_s}{\rho_a}}=246\text{ rpm}."